# LeTO: Learning Constrained Visuomotor Policy with Differentiable Trajectory Optimization

### v2

@@ V-B Arrange-chopsticks
The objective of the “Arrange-chopsticks” task is to pick up randomly placed two chopsticks and arrange them neatly on a bowl. The policy inputs end-effector poses and images from two cameras, a third-person perspective camera and a wrist-mounted camera, and controls the end-effector position, yaw angle, and grasping.
The challenges of this task include: 1. It is a relative long-horizon task, requiring the policy to have the capability to represent complex tasks. 2. Due to the slender nature of chopsticks, there is a high risk of the robot making forceful contact with the table when attempting to grasp them. For the sake of hardware safety, such collisions should be minimized as much as possible.
@@
The results are shown in Table VII. Both LeTO and diffusion policy have high success rate, but the contact forces in LeTO tests are significantly smaller.
@@ VI Discussion and future work
VI-A Hyperparameter, Training Time, and Training Stability
TABLE VIII: The training time for 20 epochs on the square ph task with a batch size of 64, conducted on a single GeForce RTX 3080. The hyperparameters for diffusion policy and LSTM-GMM are adopted from the default parameters reported in [12, 2], respectively. For the hyperparameters of LeTO, please refer to our open-source code.
@@ VI-A Hyperparameter, Training Time, and Training Stability
Regarding LeTO, its core neural network comprises CNN+LSTM, which is identical to the architecture of the basic imitation learning method BC-RNN. Therefore, we believe the discussion of hyperparameters of LeTO’s core network can refer to the discussion for BC-RNN in [12].
Concerning the hyperparameters of the DTO layer, since it comes from model-based formulation, it is possible to intuitively summarize how to select effective hyperparameters:
1. As shown in Remark 1, the optimization problem of the DTO layer is always feasible during training, given that the positions (and orientations) in the dataset meet the set position constraints. The consistent feasibility of the DTO layer ensures training stability, which is evident from the training logs of LeTO available in our open-source code.
2. In terms of constraints for acceleration and velocity, it is important to avoid overly small constraints. This could lead to a significant discrepancy between the generated trajectory and the dataset’s trajectory, adversely affecting model performance even with stable training. In our simulation experiments, LeTO trained at a maximum acceleration of 0.1 (normalized value) shows high performance, but setting it to 0.01 (normalized value) may result in worse outcomes. By analogy, the same consideration applies to the smoothing weight.
Table VIII compares the training times of LSTM-GMM, diffusion policy, and LeTO. We acknowledge that one limitation of LeTO is the computational overhead introduced by the differentiable optimization layer. Future work could look into improving the computational efficiency of the differentiable optimization process, potentially through the use of more efficient solvers or by learning approximations of the optimization layer.
@@ VI-B LeTO in Low-data and High-data Regimes
We would like to highlight that networks with the training objective of minimizing the errors for fitting human demonstrated actions can readily incorporate our proposed DTO layer at the head. This suggests that for scalability, one could consider designing a LeTO variant with transformer, enhancing the model’s capability to scale to large datasets. However, the time and cost involved in training could still be a limiting factor. Therefore, future research focused on accelerating differentiable optimization is pivotal for scalability.
@@ VI-C LeTO in Reinforcement Learning
Due to the differentiability, the DTO layer can be designed to integrated into the robot learning pipelines, enabling end-to-end optimization and inference. This opens up the possibility of extending LeTO to reinforcement learning, offering a promising future research direction. By integrating differentiable trajectory optimization with reinforcement learning, it is possible to ensure that the generated actions comply with the designed constraints, thereby significantly enhancing the safety and interpretability of the policy.
@@ VI-D LeTO in High-safety Scenarios
The safe and constraint-controlled output form of LeTO holds significant potential in robot learning scenarios where safety is of utmost importance, such as in human-robot interaction and surgical robotics. Investigating the application of LeTO, or more broadly, robot policies that integrate differential optimization in these settings, represents a highly interesting and promising research direction.

### v3

@@ V-B Arrange-chopsticks
The objective of the “Arrange-chopsticks” task is to pick up two randomly placed chopsticks and arrange them neatly on a bowl. The policy inputs end-effector poses and images from two cameras, a third-person perspective camera and a wrist-mounted camera, and controls the end-effector position, yaw angle, and grasping.
The challenges of this task include: 1. It is a relatively long-range task that requires the policy to have the capability to represent complex tasks. 2. Due to the slender nature of chopsticks, there is a high risk of the robot making forceful contact with the table when attempting to grasp them. For the sake of hardware safety, such collisions should be minimized as much as possible.
@@
The results are shown in Table X. Both LeTO and diffusion policy have high success rate, but the contact forces in LeTO tests are significantly smaller. Collisions with the environment are quite common for manipulation tasks like picking up very fine chopsticks from the tabletop. One can imagine that if the table were made of glass, such collisions could be very hazardous. Moreover, consider a scenario in which a robot frequently collides with its environment without constraints. Even if this does not affect performance in the short term, it could significantly reduce the robot’s lifespan over time.
The reduction in contact force with the tabletop is a result of the implementation of the LeTO differentiable optimization layer, which incorporates position constraints that allow for trajectory optimization. In contrast, existing ”black box” approaches, which are solely based on neural networks, do not ensure outputs that consistently meet these constraints, leading to unpredictable and potentially large impact forces with the tabletop.
Given that robots physically interact with their environment and humans, factors like safety, interpretability, stability, and robustness are crucial. Fully neural network-based approaches pose challenges for real-world deployment due to their lack of interpretability. For instance, even if the algorithm performs well under most conditions, the absence of constraints means that significant deviations in a few edge cases can lead to severe problems due to the physical nature of robots.
@@ VI Discussion and Future Work
VI-A Training Time and Training Stability
TABLE XI: The training time for 20 epochs on the square ph task with a batch size of 64, conducted on a single GeForce RTX 3080. The hyperparameters for diffusion policy and LSTM-GMM are adopted from the default parameters reported in [29, 15], respectively.
@@ VI-A Training Time and Training Stability
As shown in Remark 1, the optimization problem of the DTO layer is always feasible during training, given that the positions (and orientations) in the dataset meet the set position constraints. The consistent feasibility of the DTO layer ensures training stability, which is evident from the training logs of LeTO available in our open-source code.
Table XI compares the training times of LSTM-GMM, diffusion policy, and LeTO. We acknowledge that one limitation of LeTO is the computational overhead introduced by the differentiable optimization layer. Future work could look into improving the computational efficiency of the differentiable optimization process, potentially through the use of more efficient solvers or by learning approximations of the optimization layer.
@@ VI-B LeTO in Low-data and High-data Regimes
We would like to highlight that networks with the training objective of minimizing errors to fit human demonstrated actions can easily incorporate our proposed DTO layer at the head. This suggests that for scalability, one could consider designing a LeTO variant with transformer, enhancing the model’s capability to scale to large datasets. However, the time and cost involved in training could still be a limiting factor. As shown in Table XI, LeTO exhibits a significantly slower training compared to other methods. As highlighted in [44], solving optimization problems with precision, as done in this case, demonstrates cubic complexity relative to the number of variables and/or constraints. Therefore, future research focused on accelerating differentiable optimization is pivotal for scalability.
@@ VI-C LeTO in Reinforcement Learning
Due to the differentiability, the DTO layer can be designed to integrate into the robot learning pipelines, enabling end-to-end optimization and inference. This opens up the possibility of extending LeTO to reinforcement learning, offering a promising future research direction. By integrating differentiable trajectory optimization with reinforcement learning, it is possible to ensure that the generated actions comply with the designed constraints, thus significantly improving the safety and interpretability of the policy.
@@ VI-D LeTO in High-safety Scenarios
The safe and constraint-controlled output form of LeTO holds significant potential in robot learning scenarios where safety is of utmost importance, such as in human-robot interaction and surgical robotics. Investigating the application of LeTO, or more broadly, robot policies that integrate differentiable optimization in these settings, represents a highly interesting and promising research direction.
Additionally, since imitation learning provides a high-level controller responsible for decision-making and high-level action generation, LeTO’s trajectory optimization capabilities enable it to generate reference motions that are dynamically feasible for a low-level controller. In scenarios requiring high dynamics or heavy loads for manipulation, LeTO may have certain advantages and safety assurances when combined with more specialized lower-level controllers to accomplish these tasks.
VI-E Complex Obstacle Avoidance and Nonconvex Optimization
As stated in our paper, LeTO’s constraints must be convex. In complex environments filled with obstacles, these constraints become highly complicated and nonconvex. Therefore, LeTO does not have the ability to generate collision-free actions in confined environments filled with diverse and complex obstacles. We believe that developing an imitation learning algorithm capable of achieving collision-free performance in complex environments while executing dexterous manipulation tasks will be a very promising and important research direction.
VII acknowledgements
This work was partially supported by the United States Department of Agriculture (USDA; No. 2023-67021-39072 and 2024-67021-42878) as well as National Science Foundation (NSF; No. 2423068). This article solely reflects the opinions and conclusions of its authors and not USDA or NSF.